A 1480 kg satellite orbits the Earth at an altitude of 61.6 km. It is desired to increase the altitude of the orbit to 285 km. How much energy must be added to the system to effect this change in altitude? The universal gravitational constant is 6.67259 × 10^−11 N · m2/kg2 , the radius of the Earth is 6.37×10^6 m, and its mass is 5.98×10^24 kg.
Answer #1

The change in work done on the satellite gives the energy to be added to the system.

i.e \DeltaW = \DeltaF. \Deltax

   \DeltaW = change in work done , \DeltaF = change inforce of attraction between the satellite and the earth

   \Deltax = change in the altitudes = hf - hi = 285000 m - 61600 m

   = 223400 m

G = universal gravitational constant = 6.67259 x 10-11 Nm2/Kg2

ms = mass of satellite = 1480 Kg , Me = mass of the earth = 5.98 x1024 Kg

ri = initial distance between the centres of the satellite and the earth

= ( R + hi ) here R = radius of the earth=6.37 x106 m , hi = initial altitude = 61.6 Km = 61600 m

  = ( 6370000 m + 61600 m )

   = 6431600 m

rf = final distance between the centres of the satellite and the earth

   = ( R + hf ) here hf = final altitude = 285 Km = 285000 m

   = ( 6370000 m + 285000 m )

   = 6655000 m

Now the work done = the change in kinetic energy

  \Delta W = G ms Me [ 1/rf2   - 1/ri2 ] \Deltax

   = 6.67 x10-11 x 1480 x5.98 x1024 [ 1 / (6655000)2   - 1/ ( 6431600 )2  ] 223400

   = - 2.0685343542 x108 J here '-' indicates against to the gravity

   = 2.07 x108 J ( approx. )

So, 2.07 x108 J energy should be added to the system .